I keep getting this wrong

mitted Jan 12 at 9pm Question 7 0/0.5 pts The police department in your city was asked by the mayors office to estimate the cost of crime. The police began their study with burglary records, taking a random sample of 500 files since there were too many crime records to calculate statistics for all of the crimes committed. If the average dollar loss in a burglary, for this sample size 500 is $678, with a standard deviation of $560, construct the 95% confidence interval for the true average dollar loss in burglaries rounded to the nearest dollar amount. We can be 95% confident that the upper limit of our interval of dollar loss in crimes committed is $ 727.08 12
Answer #1

Solution:-

95% confident that the upper limit of our interval of dollar loss in crimes committed is $ 727.09.


C.I = \bar{x}\pm t_{\alpha /2}\ast \frac{\sigma }{\sqrt{n}}

C.I = 678 + 1.96*25.04396

C.I = 678 + 49.0862

C.I = (628.914, 727.086)